1.Let the gradient of a line be 3, find the equation of the line at point(1,2)
a.2y=x
b.y=2x
c.y=x
d.y=3x

2.Find the equation of the line which is parallel to
3x−2y=5 and passes through the point (-2,-4)
a.3x+2y=−5
b.3x+2y=−2
c.(-3x+2y=-2)
d.(3x+2y=6)
1
Expert's answer
2019-04-09T12:43:54-0400
  1. y-2=3(x-1) -> y=3x-1
  2. y+4=3/2(x+2) -> y=3/2x-1 -> -3x+2y=-2. Var. c.